# NOLO: Navigate Only Look Once

### v1

@@ D.2 Habitat Video Navigation Performance
2azQ1b91cZZ 50.26$\pm$ 3.10 28.99$\pm$ 1.40 311.44$\pm$ 11.24 4.30$\pm$ 0.12
8194nk5LbLH 43.94$\pm$ 1.55 22.60$\pm$ 0.51 341.28$\pm$ 1.28 4.60$\pm$ 0.08
EU6Fwq7SyZv 42.33$\pm$ 1.70 18.15$\pm$ 1.13 362.99$\pm$ 5.59 4.40$\pm$ 0.06
QUCTc6BB5sX 44.67$\pm$ 6.24 21.94$\pm$ 1.36 344.75$\pm$ 21.03 5.75$\pm$ 0.84
SN83YJsR3w2 38.33$\pm$ 6.24 15.74$\pm$ 2.48 379.77$\pm$ 17.81 4.47$\pm$ 0.97
TbHJrupSAjP 36.67$\pm$ 10.19 14.29$\pm$ 3.23 376.57$\pm$ 28.90 5.35$\pm$ 0.79
UwV83HsGsw3 30.48$\pm$ 3.75 13.66$\pm$ 0.67 402.48$\pm$ 14.42 6.30$\pm$ 0.58
Vt2qJdWjCF2 34.58$\pm$ 4.60 12.75$\pm$ 4.26 387.33$\pm$ 21.56 8.13$\pm$ 0.81
VzqfbhrpDEA 33.70$\pm$ 1.89 17.94$\pm$ 1.15 377.91$\pm$ 7.07 6.75$\pm$ 0.22
X7HyMhZNoso 66.67$\pm$ 4.38 30.84$\pm$ 2.47 275.02$\pm$ 20.74 2.87$\pm$ 0.22
Z6MFQCViBuw 40.67$\pm$ 6.18 16.25$\pm$ 3.41 369.67$\pm$ 21.35 6.12$\pm$ 0.64
gTV8FGcVJC9 53.67$\pm$ 4.03 29.49$\pm$ 2.20 292.08$\pm$ 12.03 5.48$\pm$ 0.12
gYvKGZ5eRqb 25.00$\pm$ 8.16 7.34$\pm$ 4.04 433.63$\pm$ 30.22 7.56$\pm$ 0.59
oLBMNvg9in8 45.67$\pm$ 4.03 24.45$\pm$ 3.33 335.14$\pm$ 14.15 3.69$\pm$ 0.16
pLe4wQe7qrG 70.00$\pm$ 2.72 38.63$\pm$ 2.66 218.44$\pm$ 6.09 2.38$\pm$ 0.24
pa4otMbVnkk 41.11$\pm$ 11.00 20.99$\pm$ 6.68 346.89$\pm$ 36.77 7.45$\pm$ 1.17
rqfALeAoiTq 47.50$\pm$ 3.12 21.92$\pm$ 2.23 338.93$\pm$ 11.95 4.07$\pm$ 0.11
x8F5xyUWy9e 51.82$\pm$ 0.74 23.36$\pm$ 0.97 320.58$\pm$ 4.80 4.41$\pm$ 0.06
yqstnuAEVhm 38.89$\pm$ 0.79 17.68$\pm$ 1.86 369.21$\pm$ 3.18 4.04$\pm$ 0.12
zsNo4HB9uLZ 37.08$\pm$ 3.28 18.48$\pm$ 2.66 367.31$\pm$ 11.94 5.27$\pm$ 0.34
Average 43.65 20.77 347.57 5.17
Table 12: Random Performance on Habitat Testing Set
@@
Table 13: GPT-4o Performance on Habitat Testing Set
@@
Table 14: Video-LLaVA Performance on Habitat Testing Set
@@
Table 15: NOLO-C Performance on Habitat Testing Set
@@
2azQ1b91cZZ 37.18$\pm$ 4.28 22.84$\pm$ 3.63 359.71$\pm$ 17.26 6.38$\pm$ 0.69
8194nk5LbLH 36.97$\pm$ 1.55 19.47$\pm$ 0.99 366.57$\pm$ 4.66 4.97$\pm$ 0.34
EU6Fwq7SyZv 34.00$\pm$ 2.16 18.48$\pm$ 0.86 378.66$\pm$ 2.54 5.18$\pm$ 0.31
QUCTc6BB5sX 40.33$\pm$ 1.25 23.52$\pm$ 0.77 355.46$\pm$ 1.65 6.83$\pm$ 0.32
SN83YJsR3w2 15.00$\pm$ 7.07 6.92$\pm$ 2.91 455.47$\pm$ 15.62 6.56$\pm$ 1.16
TbHJrupSAjP 20.95$\pm$ 5.99 6.05$\pm$ 1.97 437.43$\pm$ 14.11 6.53$\pm$ 0.51
UwV83HsGsw3 26.67$\pm$ 1.35 17.00$\pm$ 1.76 401.74$\pm$ 3.26 6.77$\pm$ 0.41
Vt2qJdWjCF2 19.17$\pm$ 5.62 8.34$\pm$ 0.78 437.86$\pm$ 13.29 13.49$\pm$ 0.85
VzqfbhrpDEA 25.19$\pm$ 1.05 15.50$\pm$ 2.16 398.60$\pm$ 9.64 7.91$\pm$ 1.16
X7HyMhZNoso 59.05$\pm$ 2.94 30.08$\pm$ 0.75 290.51$\pm$ 7.25 4.00$\pm$ 0.27
Z6MFQCViBuw 38.00$\pm$ 5.66 15.43$\pm$ 1.57 375.89$\pm$ 16.03 7.19$\pm$ 0.22
gTV8FGcVJC9 41.33$\pm$ 1.70 25.04$\pm$ 1.93 335.27$\pm$ 13.56 6.63$\pm$ 0.73
gYvKGZ5eRqb 11.67$\pm$ 2.36 3.51$\pm$ 1.45 474.47$\pm$ 10.70 12.91$\pm$ 0.46
oLBMNvg9in8 44.33$\pm$ 2.36 25.58$\pm$ 1.67 332.98$\pm$ 8.66 3.85$\pm$ 0.12
pLe4wQe7qrG 52.22$\pm$ 3.14 26.99$\pm$ 1.25 316.10$\pm$ 10.12 3.54$\pm$ 0.48
pa4otMbVnkk 14.44$\pm$ 5.67 7.67$\pm$ 3.04 447.38$\pm$ 20.68 9.15$\pm$ 1.53
rqfALeAoiTq 36.94$\pm$ 4.63 18.40$\pm$ 1.58 370.27$\pm$ 11.73 4.45$\pm$ 0.07
x8F5xyUWy9e 46.97$\pm$ 4.71 21.66$\pm$ 1.96 335.51$\pm$ 11.35 4.43$\pm$ 0.57
yqstnuAEVhm 38.61$\pm$ 6.17 17.60$\pm$ 3.23 372.49$\pm$ 22.68 4.89$\pm$ 0.55
zsNo4HB9uLZ 32.50$\pm$ 3.54 18.14$\pm$ 0.90 382.85$\pm$ 11.49 5.92$\pm$ 0.54
Average 33.58 17.41 381.26 6.58
Table 16: NOLO-T Performance on Habitat Testing Set
@@
2azQ1b91cZZ 41.28$\pm$ 2.97 25.50$\pm$ 2.75 344.03$\pm$ 15.14 6.45$\pm$ 0.08
8194nk5LbLH 39.70$\pm$ 2.61 19.67$\pm$ 0.74 364.86$\pm$ 3.70 5.04$\pm$ 0.27
EU6Fwq7SyZv 41.00$\pm$ 3.27 21.33$\pm$ 2.90 357.03$\pm$ 11.76 5.37$\pm$ 0.17
QUCTc6BB5sX 41.50$\pm$ 4.50 21.52$\pm$ 2.39 357.77$\pm$ 11.40 7.51$\pm$ 0.73
SN83YJsR3w2 21.67$\pm$ 11.79 7.83$\pm$ 4.20 432.92$\pm$ 32.98 7.79$\pm$ 1.89
TbHJrupSAjP 29.52$\pm$ 0.67 11.53$\pm$ 0.91 400.50$\pm$ 5.36 6.45$\pm$ 0.39
UwV83HsGsw3 28.10$\pm$ 0.67 15.16$\pm$ 2.66 398.37$\pm$ 9.11 6.95$\pm$ 0.19
Vt2qJdWjCF2 18.33$\pm$ 3.12 7.74$\pm$ 1.46 443.81$\pm$ 10.96 13.66$\pm$ 0.67
VzqfbhrpDEA 30.00$\pm$ 0.91 16.90$\pm$ 0.92 387.73$\pm$ 5.64 9.28$\pm$ 0.43
X7HyMhZNoso 63.33$\pm$ 0.89 32.24$\pm$ 0.69 270.01$\pm$ 2.15 4.31$\pm$ 0.39
Z6MFQCViBuw 40.67$\pm$ 3.40 15.79$\pm$ 3.40 380.14$\pm$ 11.92 6.87$\pm$ 0.44
gTV8FGcVJC9 46.50$\pm$ 1.50 28.97$\pm$ 1.96 313.77$\pm$ 0.72 7.54$\pm$ 0.94
gYvKGZ5eRqb 10.00 2.73$\pm$ 0.50 471.45$\pm$ 3.05 11.34$\pm$ 1.00
oLBMNvg9in8 47.00$\pm$ 1.00 24.69$\pm$ 1.35 332.24$\pm$ 4.97 3.92$\pm$ 0.16
pLe4wQe7qrG 65.00$\pm$ 1.67 28.04$\pm$ 5.24 290.87$\pm$ 27.90 2.58$\pm$ 0.09
pa4otMbVnkk 33.33$\pm$ 3.33 18.79$\pm$ 0.48 365.55$\pm$ 13.95 9.17$\pm$ 0.66
rqfALeAoiTq 43.89$\pm$ 2.19 22.01$\pm$ 1.75 348.34$\pm$ 8.51 4.55$\pm$ 0.29
x8F5xyUWy9e 52.42$\pm$ 0.43 22.44$\pm$ 0.94 328.36$\pm$ 2.28 4.13$\pm$ 0.02
yqstnuAEVhm 38.61$\pm$ 3.14 17.06$\pm$ 2.82 372.31$\pm$ 14.27 4.83$\pm$ 0.13
zsNo4HB9uLZ 26.67$\pm$ 1.18 16.10$\pm$ 1.78 400.21$\pm$ 5.77 6.74$\pm$ 0.08
Average 37.93 18.80 368.01 6.72
Table 17: NOLO(M) Performance on Habitat Testing Set
@@
2azQ1b91cZZ 35.77$\pm$ 0.38 21.70$\pm$ 0.98 367.77$\pm$ 0.44 7.26$\pm$ 0.06
8194nk5LbLH 38.18$\pm$ 0.91 19.20$\pm$ 0.74 365.92$\pm$ 3.05 5.05$\pm$ 0.18
EU6Fwq7SyZv 40.50$\pm$ 4.50 20.97$\pm$ 1.42 360.27$\pm$ 14.82 5.04$\pm$ 0.49
QUCTc6BB5sX 40.00 22.11$\pm$ 1.36 354.65$\pm$ 1.51 6.64$\pm$ 0.25
SN83YJsR3w2 32.50$\pm$ 7.50 10.28$\pm$ 2.55 414.92$\pm$ 23.22 6.92$\pm$ 1.20
TbHJrupSAjP 24.29$\pm$ 1.43 7.48$\pm$ 0.50 427.30$\pm$ 6.51 6.63$\pm$ 0.07
UwV83HsGsw3 30.71$\pm$ 2.14 17.89$\pm$ 3.05 390.06$\pm$ 15.81 6.03$\pm$ 0.21
Vt2qJdWjCF2 13.75$\pm$ 1.25 6.65$\pm$ 0.62 454.84$\pm$ 0.97 14.59$\pm$ 0.18
VzqfbhrpDEA 28.33$\pm$ 3.89 14.89$\pm$ 2.38 402.32$\pm$ 9.62 9.08$\pm$ 0.47
X7HyMhZNoso 64.64$\pm$ 1.79 30.95$\pm$ 0.11 276.48$\pm$ 5.50 3.94$\pm$ 0.15
Z6MFQCViBuw 37.00$\pm$ 7.00 16.86$\pm$ 3.37 374.19$\pm$ 22.49 6.90$\pm$ 0.33
gTV8FGcVJC9 50.50$\pm$ 7.50 28.67$\pm$ 1.23 306.25$\pm$ 9.88 6.54$\pm$ 0.72
gYvKGZ5eRqb 12.50$\pm$ 2.50 3.79$\pm$ 1.73 475.62$\pm$ 14.77 11.55$\pm$ 1.20
oLBMNvg9in8 48.00$\pm$ 2.00 28.67$\pm$ 0.73 319.90$\pm$ 6.13 3.61$\pm$ 0.16
pLe4wQe7qrG 53.33$\pm$ 3.33 26.82$\pm$ 1.01 311.75$\pm$ 0.38 3.26$\pm$ 0.19
pa4otMbVnkk 20.00 8.13$\pm$ 0.40 431.58$\pm$ 1.42 10.87$\pm$ 0.33
rqfALeAoiTq 45.42$\pm$ 2.08 20.42$\pm$ 2.06 352.25$\pm$ 7.54 4.19$\pm$ 0.23
x8F5xyUWy9e 49.55$\pm$ 4.09 22.50$\pm$ 1.80 331.82$\pm$ 12.30 4.14$\pm$ 0.47
yqstnuAEVhm 42.08$\pm$ 4.58 18.98$\pm$ 3.12 360.02$\pm$ 16.95 4.89$\pm$ 0.05
zsNo4HB9uLZ 27.50 16.36$\pm$ 1.47 400.51$\pm$ 1.49 6.76$\pm$ 0.10
Average 36.73 18.17 373.92 6.69
D.3 Cross Evaluation Results
Table 18: NOLO Performance in RoboTHOR after training in Habitat

### v2

@@ 9.2 Habitat Evaluation Performance
2azQ1b91cZZ 50.26$\pm$ 3.10 28.99$\pm$ 1.40 311.44$\pm$ 11.24 2.80$\pm$ 0.12
8194nk5LbLH 43.94$\pm$ 1.55 22.60$\pm$ 0.51 341.28$\pm$ 1.28 3.10$\pm$ 0.08
EU6Fwq7SyZv 42.33$\pm$ 1.70 18.15$\pm$ 1.13 362.99$\pm$ 5.59 2.90$\pm$ 0.06
QUCTc6BB5sX 44.67$\pm$ 6.24 21.94$\pm$ 1.36 344.75$\pm$ 21.03 4.25$\pm$ 0.84
SN83YJsR3w2 38.33$\pm$ 6.24 15.74$\pm$ 2.48 379.77$\pm$ 17.81 2.97$\pm$ 0.97
TbHJrupSAjP 36.67$\pm$ 10.19 14.29$\pm$ 3.23 376.57$\pm$ 28.90 3.85$\pm$ 0.79
UwV83HsGsw3 30.48$\pm$ 3.75 13.66$\pm$ 0.67 402.48$\pm$ 14.42 4.80$\pm$ 0.58
Vt2qJdWjCF2 34.58$\pm$ 4.60 12.75$\pm$ 4.26 387.33$\pm$ 21.56 6.63$\pm$ 0.81
VzqfbhrpDEA 33.70$\pm$ 1.89 17.94$\pm$ 1.15 377.91$\pm$ 7.07 5.25$\pm$ 0.22
X7HyMhZNoso 66.67$\pm$ 4.38 30.84$\pm$ 2.47 275.02$\pm$ 20.74 1.37$\pm$ 0.22
Z6MFQCViBuw 40.67$\pm$ 6.18 16.25$\pm$ 3.41 369.67$\pm$ 21.35 4.62$\pm$ 0.64
gTV8FGcVJC9 53.67$\pm$ 4.03 29.49$\pm$ 2.20 292.08$\pm$ 12.03 3.98$\pm$ 0.12
gYvKGZ5eRqb 25.00$\pm$ 8.16 7.34$\pm$ 4.04 433.63$\pm$ 30.22 6.06$\pm$ 0.59
oLBMNvg9in8 45.67$\pm$ 4.03 24.45$\pm$ 3.33 335.14$\pm$ 14.15 2.19$\pm$ 0.16
pLe4wQe7qrG 70.00$\pm$ 2.72 38.63$\pm$ 2.66 218.44$\pm$ 6.09 0.88$\pm$ 0.24
pa4otMbVnkk 41.11$\pm$ 11.00 20.99$\pm$ 6.68 346.89$\pm$ 36.77 5.95$\pm$ 1.17
rqfALeAoiTq 47.50$\pm$ 3.12 21.92$\pm$ 2.23 338.93$\pm$ 11.95 2.57$\pm$ 0.11
x8F5xyUWy9e 51.82$\pm$ 0.74 23.36$\pm$ 0.97 320.58$\pm$ 4.80 2.91$\pm$ 0.06
yqstnuAEVhm 38.89$\pm$ 0.79 17.68$\pm$ 1.86 369.21$\pm$ 3.18 2.54$\pm$ 0.12
zsNo4HB9uLZ 37.08$\pm$ 3.28 18.48$\pm$ 2.66 367.31$\pm$ 11.94 3.77$\pm$ 0.34
Average 43.65 20.77 347.57 3.67
Table 16: Random performance in Habitat testing scenes.
@@
Table 17: GPT-4o performance in Habitat testing scenes.
@@
Table 18: Video-LLaVA performance in Habitat testing scenes.
@@
Table 19: AVDC performance in Habitat testing scenes.
@@
2azQ1b91cZZ 24.12 7.93 429.30 3.44
8194nk5LbLH 29.33 14.02 375.06 5.05
EU6Fwq7SyZv 26.95 12.35 442.17 5.45
QUCTc6BB5sX 22.14 10.48 450.81 3.54
SN83YJsR3w2 30.77 15.06 395.85 4.21
TbHJrupSAjP 31.49 10.74 420.21 4.96
UwV83HsGsw3 23.57 12.84 393.14 3.33
Vt2qJdWjCF2 28.05 11.54 426.06 4.42
VzqfbhrpDEA 25.22 14.19 410.55 3.45
X7HyMhZNoso 29.58 12.82 414.27 4.20
Z6MFQCViBuw 26.78 13.78 430.31 4.63
gTV8FGcVJC9 28.29 10.00 411.56 4.27
gYvKGZ5eRqb 26.52 10.18 421.96 4.92
oLBMNvg9in8 25.44 13.00 417.07 3.91
pLe4wQe7qrG 27.88 16.95 429.00 4.40
pa4otMbVnkk 29.31 10.21 434.91 5.12
rqfALeAoiTq 28.45 12.56 403.33 4.42
x8F5xyUWy9e 24.68 9.76 432.66 3.68
yqstnuAEVhm 30.19 13.00 399.91 4.39
zsNo4HB9uLZ 26.94 13.32 424.30 4.14
Average 27.29 12.24 418.12 4.30
Table 20: VGM performance in Habitat testing scenes.
@@
2azQ1b91cZZ 35.88 16.12 373.11 4.48
8194nk5LbLH 31.27 19.02 423.51 3.85
EU6Fwq7SyZv 34.10 18.74 368.62 4.06
QUCTc6BB5sX 36.27 16.88 367.95 4.29
SN83YJsR3w2 32.92 17.09 426.29 4.73
TbHJrupSAjP 35.08 18.43 395.03 4.40
UwV83HsGsw3 33.45 16.21 424.39 3.82
Vt2qJdWjCF2 36.12 15.91 403.36 4.64
VzqfbhrpDEA 33.57 19.50 397.14 3.67
X7HyMhZNoso 34.88 18.62 398.75 4.05
Z6MFQCViBuw 32.67 15.53 398.10 4.37
gTV8FGcVJC9 35.13 17.68 390.74 4.36
gYvKGZ5eRqb 34.77 16.44 393.34 4.36
oLBMNvg9in8 31.22 17.13 391.16 3.38
pLe4wQe7qrG 36.45 18.04 357.49 4.19
pa4otMbVnkk 34.02 16.66 414.59 3.99
rqfALeAoiTq 35.23 18.21 369.25 4.20
x8F5xyUWy9e 32.45 17.90 380.19 4.32
yqstnuAEVhm 34.99 15.80 412.49 4.17
zsNo4HB9uLZ 36.23 16.09 392.92 4.39
Average 34.34 17.30 393.92 4.19
Table 21: ZSON performance in Habitat testing scenes.
@@
2azQ1b91cZZ 30.22 21.90 400.05 4.49
8194nk5LbLH 35.14 17.42 397.11 4.13
EU6Fwq7SyZv 31.88 18.15 360.23 4.71
QUCTc6BB5sX 34.62 21.22 379.75 4.42
SN83YJsR3w2 30.87 19.88 415.70 4.21
TbHJrupSAjP 33.94 18.75 428.33 4.05
UwV83HsGsw3 35.56 20.32 391.88 3.92
Vt2qJdWjCF2 33.18 22.58 400.38 4.37
VzqfbhrpDEA 29.34 19.49 393.52 4.72
X7HyMhZNoso 34.90 18.22 376.97 4.29
Z6MFQCViBuw 32.12 21.74 380.55 4.91
gTV8FGcVJC9 30.25 17.96 406.00 4.35
gYvKGZ5eRqb 35.44 22.08 412.43 3.81
oLBMNvg9in8 31.53 19.14 366.82 4.58
pLe4wQe7qrG 30.87 20.69 393.75 4.14
pa4otMbVnkk 33.97 18.56 418.57 4.28
rqfALeAoiTq 34.12 19.24 378.22 4.33
x8F5xyUWy9e 32.78 18.97 416.39 3.91
yqstnuAEVhm 31.41 21.66 419.74 4.27
zsNo4HB9uLZ 34.55 19.31 382.41 4.43
Average 32.83 19.86 395.94 4.32
Table 22: NOLO-C performance in Habitat testing scenes.